# Off-Equatorial Orbits around Magnetically Charged Black Holes

### v1

@@ IV Orbital structure in Rotating MBH
As in Section II, we define the electromagnetic potential on the northern-hemisphere patch, while the southern patch is related by the gauge transformation. To derive the electromagnetic tensor in the rotating magnetically charged spacetime, we follow the complex transformation algorithm [Newman1965, Adamo2014], starting from the static spherically symmetric MBH described by the Reissner–Nordström metric.
@@
$A_{\mu}=\left(\frac{-Pa\cos\theta}{\rho^{2}},0,0,\frac{r^{2}+a^{2}}{\rho^{2}}P\cos\theta\right).$ (58)
@@
$F_{\mu\nu}=\begin{bmatrix}0&-\frac{2Par\cdot\cos\theta}{\rho^{4}}&-\frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\\ \frac{2Par\cdot\cos\theta}{\rho^{4}}&0&0&-\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}\\ \frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0&0&-\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}\\ 0&\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}&\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\end{bmatrix}$. (59)
In a rotating charged spacetime, the behavior of particles differs between the prograde and retrograde cases. Following a similar argument as in Section II, the photon ring radius is determined by [Johannsen2013, Gal2019]
$r_{\text{pr}}^{2}-3Mr_{\text{pr}}+2P^{2}\pm 2a\sqrt{Mr_{\text{pr}}-P^{2}}=0,$ (60)
where ++ and - denote prograde and retrograde motion, respectively. For extreme MBHs, $r_{\text{ph}}$ is identical to the horizon radius $r_{\text{h}}=M$.
The ISCO radius is obtained by defining [Liu2017]
$\begin{split}\mathcal{R}(r):=&\rho^{4}(\frac{dr}{d\tau})^{2}\\ =&\pm 2aj\varepsilon(P^{2}-2Mr)-(\Delta-a^{2})(r^{2}+j^{2})\\ &+r^{4}\varepsilon^{2}+a^{2}\left(\varepsilon^{2}(r^{2}+2Mr-P^{2})-r^{2}\right),\end{split}$ (61)
where ++ and - denote prograde and retrograde motion, $\varepsilon$ and $j$ denote the energy and angular momentum per unit mass. Stable circular orbits satisfy
$\mathcal{R}(r)=\frac{d\mathcal{R}(r)}{dr}=0,\quad\frac{d^{2}\mathcal{R}(r)}{dr^{2}}\geq 0,$ (62)
which yields the ISCO condition
$\begin{split}Mr(6Mr-r^{2}-9P^{2}+3a^{2})&+4P^{2}(P^{2}-a^{2})\\ \mp 8a(Mr-P^{2})^{3/2}&=0.\end{split}$ (63)
Here - and ++ denote prograde and retrograde motion, respectively.
The electromagnetic field in this spacetime diverts electrically charged particles from the equator in a similar way. While the electromagnetic field in this spacetime is complicated, we again have off-equatorial circular orbits similar to those discussed in Section II.
We begin by examining the $t$ and $\varphi$ components of the geodesic equations Eq. (6). It turns out that the relationships in Eq. (13) still hold for a rotating MBH. Consequently, for circular orbits with $u^{r}=u^{\theta}=0$, all terms in these equations vanish.
Figure 5: $\theta(r)$ of a q/m=1q/m=-1 particle prograding a $P/M=0.6$ spinning MBH, for multiple $P/M$ values. The peak and smallest allowed radius decrease monotonously with $a/M$.
The remaining two equations determine the angular velocity and polar angle of a particle on a spherical orbit. Due to their complexity, the two variables, $\theta$ and $\omega$, are coupled in the equations of motion. Their explicit forms are relegated to Appendix A and can only be solved numerically. It is worth noting that the Carter constant can be used to decouple the radial and polar equations [Dyson2023]. Nevertheless, this approach introduces offsetting complexities into the temporal and azimuthal components of the motion. We therefore proceed with the direct numerical solution of the coupled system.
Because of the periodicity of the trigonometric functions involved, these equations yield infinitely many solutions. However, within the range $0\leq\theta\leq\pi$, only two distinct branches exist, corresponding to prograde and retrograde particles. These two branches are shown in Figure 5 (prograde) and Figure 6 (retrograde), respectively. In each plot, we fix the particle charge-to-mass ratio at q/m=1q/m=-1, set the magnetic charge to $P=0.6M$, and vary the rotation parameter from $a=0$ to $a=0.8M$, approaching the extremal case.
Figure 6: $\theta(r)$ of a q/m=1q/m=-1 particle retrograding a $P/M=0.6$ spinning MBH, for multiple $P/M$ values. The peak and smallest allowed radius increase monotonously with $a/M$.
The behavior of $\theta(r)$ is qualitatively similar to the non-rotating case: it approaches $\pi/2$ in both the small- and large-radius limits, with a single intermediate peak. An exception arises for prograde motion around an extremal MBH, where $\theta$ attains its extremum near the smallest admissible radius. For retrograde orbits, both the peak and the innermost allowed radius occur farther from the black hole, consistent with the intuitive picture that prograde motion is more closely aligned with the spacetime geometry. Contrary to Section II, the ISCO is not well defined for charged particles in this spacetime, since $\mathcal{R}(r)$ depends on three constants, energy, angular momentum, and the Carter constant [Dyson2023], thus the ISCO condition (Eq. (62)) does not yield a unique marginal orbit. Consequently, the peak radius lacks a direct physical interpretation.
Despite this, the photon ring still serves as the absolute inner boundary for circular orbits. To demonstrate it, we analyze the radial component of the equation of motion, which determines the radial acceleration $R(r,\theta,\omega)=d^{2}r/d\tau^{2}$. A circular orbit can only exist if there is a set of parameters $(\theta,\omega)$ for which $R=0$. We numerically computed the maximum possible value of $R$ over all latitudes and a wide range of angular velocities at a given radius $r$.
Figure 7: Maximum radial acceleration $R(r,\theta,\omega)$ over all latitudes and $M\omega\leq 10$, as a function of the relative radius to photon ring radius $r/r_{\text{pr}}$. The black dashed lines mark $r=r_{\text{pr}}$ and $R=0$. A sharp transition occurs at the photon ring radius $r=r_{\text{pr}}$: $\max(R)<0$ for $r<r_{\text{pr}}$, prohibiting circular motion, while $\max(R)>0$ for $r>r_{\text{pr}}$, allowing circular orbits.
The results are shown in Figure 7. There is a sharp transition at the photon ring radius ($r=r_{\text{pr}}$). For all radii inside the photon ring ($r<r_{\text{pr}}$), the maximum value of $R$ is negative. This implies that no amount of orbital velocity can counteract the gravitational pull, and any particle is destined to plunge. For $r>r_{\text{pr}}$, $\max(R)$ becomes positive, allowing for solutions where $R=0$. Thus, the photon ring marks the critical inner boundary for the existence of off-equatorial circular orbits for charged particles in the rotating spacetime.
In Figure 7 we show the maximum of $R(\theta,\omega)$ in the range of $0\leq\theta\leq\pi,\quad 0\leq M\omega\leq 10$ as a function of the relative radius to photon ring radius ($r/r_{\text{pr}}$), for prograde and retrograde test particle with q/m=1q/m=-1, for $a/M$ from 0 to extremum (denoted by color from purple to yellow) at three different $P/M$. There is a jump point at $r_{\text{pr}}$: inside the photon ring radius, the maximum of $R(\theta,\omega)$ is negative, thus excluding any solution of a circular orbit; when $r>r_{\text{pr}}$ however, the maximum depends on the range of $\omega$ but tends to infinity as $\omega$ increases.
We then conclude that $R(\theta,\omega)$ has no root at $r<r_{\text{pr}}$, and from Section IV, we see that Eq. (6) always have solution. Thus, the photon ring radius marks the critical radius above which there exist off-equatorial circular orbits for charged particles.
Finally, it is worth noting that rotating electrically charged black holes also possess radial and polar components of both the electric and magnetic field. However, in contrast to the magnetically charged case, no off-equatorial circular orbits exist here. For an electrically charged particle, circular motion is confined to the equatorial plane only.
Figure 8: $\Theta(\theta,\omega,r)$ in Kerr–Newman space-time with $a/M=Q/M=0.5$, evaluated at a wide range of $\omega$, and three radii: the photon ring ($r=r_{\text{pr}}$), the ISCO, and a large-$r$ regime, for a prograde q/m=1q/m=-1 particle. At all radii, $\Theta$ vanishes only in the equatorial limit $\theta\to\pi/2$. A secondary trend is visible: near the photon ring, $\Theta$ is suppressed at large $\omega$, while at large radii it is suppressed at small $\omega$, reflecting the shift in angular motion with radius.
The electromagnetic tensor in Kerr–Newman spacetime takes the form
To show this explicitly, we define
$\Theta(\theta,\omega,r)=d^{2}\theta/d\tau^{2}\,$, (65)
as the total polar acceleration. Figure 8 plots $\Theta$ for a charged particle in an electrically charged KN spacetime. Away from the equatorial plane, $\Theta$ is strictly non-zero, confirming that no equilibrium solution exists. This highlights a fundamental difference: the purely radial field of a magnetic monopole provides a unique mechanism for latitudinal confinement that is absent in the field of a rotating electric charge.
V Conclusions
In conclusion, we have presented a systematic and detailed analysis of the orbital dynamics of charged test particles around both static and rotating magnetically charged black holes. Although the motion of neutral particles in such spacetimes is well understood, we have demonstrated that the interplay between the gravitational field and the long-range magnetic Lorentz force gives rise to a novel family of stable, off-equatorial circular orbits. These trajectories, confined to a fixed latitude, represent a significant departure from the planar motion typically associated with black hole orbits.
Our primary results were derived for the static, spherically symmetric MBH. We obtained a closed-form analytical expression, Eq. (16), which determines the equilibrium polar angle $\theta$ of a charged particle as a function of its orbital radius. Through a rigorous analysis, we established profound connections between these new orbits and the fundamental structure of the spacetime including the photon sphere and ISCO. Our analysis of synchrotron radiation confirms these orbits are viable over astrophysical timescales. We found that even electrons, which maximizes radiative losses, can maintain a stable orbit with a significant latitudinal deviation.
We extended this investigation to rotating MBH spacetime, where we numerically computed the prograde and retrograde orbital branches. Our results show that the MBH spin significantly modifies the orbits, pulling prograde trajectories closer and pushing retrograde ones farther away. Despite the added complexity and the loss of a direct ISCO connection, we demonstrated that the photon ring robustly maintains its role as the inner boundary for stable circular motion. We also highlighted a crucial distinction by showing that such off-equatorial orbits are forbidden in the spacetime of an electrically charged Kerr-Newman black hole, underscoring the unique role of the radial magnetic field in providing latitudinal confinement.
VI Acknowledgments
The authors gratefully acknowledge Julian Krolik and Emanuele Berti for their stimulating discussions. We thank David Pereñiguez for his insightful comments on the manuscript. Xilai Li would like to thank the Shanghai Jiao Tong University International Cooperation and Exchange Office and the Johns Hopkins University Global Education Office for facilitating a visit to JHU. L.D.G. is supported by NSF Grants No. PHY-2207502, AST-2307146, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, and by the Simons Investigator Grant No. 144924.
Appendix A Equations of Motion of Charged Particles Around a Rotating MBH

### v2

@@ IV.1 Circular Orbits in Rotating MBHs
The electromagnetic potential for the rotating MBH black holes is [16]
$A_{\mu}=\frac{P\cos\theta}{\rho^{2}}\left(a,0,0,-(r^{2}+a^{2})\right)\,$, (49)
which reproduces Eq. (4) in the limit $a\to 0$ up to a constant. We point the reader to Appendix A for a derivation of Eq. (49).
Following the same procedure outlined in Section II, we solve the equations of motion under the assumption of a circular orbit, $u^{\mu}=(u^{t},0,0,u^{\varphi})$, where $u^{\varphi}$ is the constant angular velocity (its sign distinguishes between prograde and retrograde motion).
Figure 5: $\theta(r)$ of a q/m=1q/m=-1 particle prograding a $P/M=0.6$ spinning MBH, for multiple $a/M$ values. Solid lines denote stable orbits, while dashed lines indicate instability against small perturbations; the transition point is marked on each curve. The y-axis is inverted so that north pole direction ($\theta=0$) appears at the top of the figure, matching the intuitive orientation. The peak and smallest allowed radius decrease monotonously with $a/M$.
Figure 6: $\theta(r)$ of a q/m=1q/m=-1 particle retrograding a $P/M=0.6$ spinning MBH, for multiple $a/M$ values. Solid lines denote stable orbits, while dashed lines indicate instability against small perturbations; the transition point is marked on each curve. The y-axis is inverted so that north pole direction ($\theta=0$) appears at the top of the figure, matching the intuitive orientation. The peak and smallest allowed radius increase monotonously with $a/M$.
The $t$-component of the four velocity is determined imposing the constraint $g_{\mu\nu}u^{\mu}u^{\nu}=-1$, whereas Eq (12) leads to two independent equations that are used to numerically determine the polar angle $\theta$ and the radius $r$ of each orbit as a function of the parameter $u^{\varphi}$ under the condition $d^{2}r/d\tau^{2}=d^{2}\theta/d\tau^{2}=0$ (see Appendix B for more details). For completeness, it is worth mentioning the Carter constant can be used to decouple the radial and polar equations [17] (see also Sec. IV.3). However, we proceed with numerically finding the roots of Eq. (66) and Eq. (67), as this approach allows us to directly recover the family of circular orbits as a function of the spacetime coordinates $(r,\theta)$ and the parameters $(P/M,a/M,q/m)$ alone. The results for the prograde and retrograde branches are shown in Figure 5 and 6, respectively. In each plot, we fix the particle charge-to-mass ratio at q/m=1q/m=-1, set the magnetic charge to $P=0.6M$, and vary the rotation parameter from $a=0$ to $a=0.8M$, which corresponds to the extremal case according to Eq. (48). Analogous to the static case, we identify the transition from stable to unstable orbits and denote unstable regions with dashed lines. As will be discussed in Section IV.2, this transition point still coincides with the peak radius of the orbital profile.
Figure 7: Maximum radial acceleration $MR(r,\theta,\omega)$ over all latitudes and $M\omega\leq 10$, as a function of the relative radius to photon ring radius $r/r_{\text{pr}}$. The black dashed lines mark $r=r_{\text{pr}}$ and $R=0$. A sharp transition occurs at the photon ring radius $r=r_{\text{pr}}$: $\max(R)<0$ for $r<r_{\text{pr}}$, prohibiting circular motion, while $\max(R)>0$ for $r>r_{\text{pr}}$, allowing circular orbits.
The behavior of $\theta(r)$ is qualitatively similar to the non-rotating case: it approaches $\pi/2$ in both the small- and large-radius limits, with a single intermediate peak. An exception arises for prograde motion around an extremal MBH, where $\theta$ attains its extremum near the smallest admissible radius. For retrograde orbits, both the peak and the innermost allowed radius occur farther from the black hole, consistent with the intuitive picture that prograde motion is more closely aligned with the spacetime geometry.
IV.2 Photon Rings and Stability Analysis
Following a similar argument as in Section II, the photon ring radius in the equatorial plane is determined by the roots of the equation [26, 19]
$r_{\text{pr}}^{2}-3Mr_{\text{pr}}+2P^{2}\pm 2a\sqrt{Mr_{\text{pr}}-P^{2}}=0,$ (50)
where ++ and - denote prograde and retrograde motion, respectively. For extreme MBHs, $r_{\text{pr}}$ is identical to the horizon radius $r_{\text{h}}=M$. Our numerical exploration reveals that the photon ring continues to serve as the absolute inner boundary for circular orbits, in analogy with the non-rotating case. To demonstrate it, we analyze the radial component of the equation of motion, which determines the radial acceleration $R(r,\theta,\omega)=d^{2}r/d\tau^{2}$. A circular orbit can only exist if there is a set of parameters $(\theta,\omega)$ for which $R=0$. We numerically computed the maximum possible value of $R$ over all latitudes and a wide range of angular velocities at a given radius $r$. The results are presented in Figure 7, where we show the maximum of $R(\theta,\omega)$ in the range of $0\leq\theta\leq\pi,\quad 0\leq M\omega\leq 10$ as a function of the relative radius to photon ring radius ($r/r_{\text{pr}}$), for prograde and retrograde test particle with q/m=1q/m=-1, for $a/M$ from 0 to extremum (denoted by color from purple to yellow) at three different $P/M$. There is a jump point at $r_{\text{pr}}$: inside the photon ring radius, the maximum of $R(\theta,\omega)$ is negative, thus excluding any solution of a circular orbit; when $r>r_{\text{pr}}$ however, the maximum depends on the range of $\omega$ but tends to infinity as $\omega$ increases.
Consequently, $R(\theta,\omega)$ has no root at $r<r_{\text{pr}}$, and from Section IV.1, we see that Eq. (12) always have solution. Thus, the photon ring radius marks the critical radius above which there exist off-equatorial circular orbits for charged particles.
Unlike the static scenario, the classical Innermost Stable Circular Orbit (ISCO) is not uniquely defined for charged particles in rotating MBH spacetimes, as the radial motion depends on a third constant of motion (the Carter constant) in addition to energy and angular momentum. Consequently, a family of marginal stability radii exists for different Carter constants. We address this by defining a stability criterion specific to the class of orbits computed in Section IV.1.
We consider a charged test particle initially on an equilibrium circular orbit at $(r_{0},\theta(r_{0}))$. We impose a small radial perturbation $r^{\prime}=r_{0}+\delta_{r}$ while constraining the particle to the manifold of circular solutions, such that the new polar angle is $\theta^{\prime}=\theta(r^{\prime})$. The perturbed angular velocity $\omega^{\prime}$ is determined via angular momentum conservation. We then evaluate the radial acceleration at the perturbed radius $r^{\prime}$ and define the stability coefficient $\kappa$ as the derivative of radial acceleration with respect to radial displacement
$\kappa(r_{0})=\frac{d^{2}r/d\tau^{2}|_{r=r_{0}+\delta_{r}}}{\delta_{r}}.$ (51)
Figure 8: Normalized absolute value of stability coefficient $\kappa$ (log scale) versus radius for a prograde particle (q/m=1q/m=-1) orbiting spinning MBHs ($P/M=0.6$). The function transitions from positive (unstable) to negative (stable), creating a sharp trough in the log plot. Dashed vertical lines indicate the peak radius of $\theta(r)$ for each spin parameter, which matches the stability transition point within numerical error.
Figure 9: Normalized absolute value of stability coefficient $\kappa$ (log scale) versus radius for a retrograde particle (q/m=1q/m=-1) orbiting spinning MBHs ($P/M=0.6$). Similar to the prograde case, the sharp trough corresponds to the sign change of the stability coefficient. The dashed vertical lines confirm that the stability transition coincides with the peak radius of the $\theta(r)$ profile.
A negative value of $\kappa$ indicates a restoring force, signaling stability against small perturbations, while a positive value implies instability. In Figures 8 and 9, we plot the magnitude of $\kappa$ in log scale for the configuration used in Section IV.1. The force is consistently positive for radii smaller than the peak radius and negative for larger radii. The peak of the $\theta(r)$ profile therefore continues to demarcate the transition from unstable to stable orbits.
We therefore conclude that the minimum allowed radius for circular orbits coincides with the photon ring radius, and that the peak radius of the $\theta(r)$ function marks the innermost stable orbit.
IV.3 The case of electrically charged rotating black holes
Finally, it is worth noting that rotating electrically charged black holes also possess radial and polar components of both the electric and magnetic field. However, in contrast to the magnetically charged case, no off-equatorial circular orbits exist here. In other words, for an electrically charged particle, circular motion is confined to the equatorial plane only.
To see this, let us recall that the electromagnetic potential for the rotating electrically charged black hole is [16]
$A_{\mu}=\frac{Qr}{\rho^{2}}\left(-1,0,0,a\sin^{2}\theta\right)\,.$ (52)
While the energy $E$ and angular momentum $L$ per unit mass are conserved due to the existence of two Killing vectors, an additional constant of motion can be identified using the Hamilton–Jacobi approach. This quantity, known as the Carter constant [10] and hereafter denoted as $\mathcal{C}$, ensures the integrability of the equations of motion. In particular, the equation for the polar angle can be written as [16]
$\left(\frac{dz}{d\tau}\right)^{2}\!\!\!=\!\frac{1}{\rho^{4}}\left\{\mathcal{C}(1-z^{2})-z^{2}[L^{2}+a^{2}(1-z^{2})(1-E^{2})]\right\}\,$, (53)
where $z=\cos\theta$. A circular orbit must satisfy $dz/d\tau=0$ and $d^{2}z/d\tau^{2}=0$. The former condition is used to express $\mathcal{C}$ in terms of the other parameters,
$\mathcal{C}=z^{2}\left(-a^{2}\left(E^{2}-1\right)-\frac{L^{2}}{z^{2}-1}\right)\,.$ (54)
Substituting Eq. (54) into the further condition $d^{2}z/d\tau^{2}=0$, and asking that the orbit must be bound, i.e. $E^{2}<1$, leads to the unique solution $z=0$. Thus, circular orbits in electrically charged rotating black holes exist only in the equatorial plane and are characterized by a vanishing Carter constant.
V Conclusions
In conclusion, we have presented a systematic and detailed analysis of the orbital dynamics of charged test particles around both static and rotating magnetically charged black holes. Although the motion of neutral particles in such spacetimes is well understood, we have demonstrated that the interplay between the gravitational field and the long-range magnetic Lorentz force gives rise to a family of stable, off-equatorial circular orbits. These trajectories, confined to a fixed latitude, represent a significant departure from the planar motion typically associated with black hole orbits.
Our primary results were derived for the static, spherically symmetric MBH. We obtained a closed-form analytical expression, Eq. (17), which determines the equilibrium polar angle $\theta$ of a charged particle as a function of its orbital radius. This relation establishes a direct link between fundamental radii: the minimum orbital radius corresponds to the photon ring, while the profile peak coincides with the ISCO. Our analysis of synchrotron radiation indicates these orbits are viable over astrophysical timescales. We found that even electrons, which maximize radiative losses, can maintain a stable orbit with a significant latitudinal deviation around astrophysical-sized black holes.
We extended this investigation to rotating MBH spacetime, where we numerically computed the prograde and retrograde orbital branches. Our results show that the MBH spin modifies the orbits, while keeping the similar shape as the static case. Despite the added complexity and the loss of a direct ISCO connection, we demonstrated that the least allowed radius and the peak radius robustly maintain their roles as the photon ring and the inner boundary for stable circular motion, respectively. We also highlighted a crucial distinction by showing that such off-equatorial orbits are forbidden in electrically charged Kerr-Newman spacetimes, underscoring the unique role of the magnetic charge.
VI Acknowledgments
The authors gratefully acknowledge Julian Krolik and Emanuele Berti for their stimulating discussions. We thank David Pereñiguez for his insightful comments on the manuscript. L.X. would like to thank the Shanghai Jiao Tong University International Cooperation and Exchange Office and the Johns Hopkins University Global Education Office for facilitating a visit to JHU. L.D.G. is supported by NSF Grants No. PHY-2207502, AST-2307146, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, and by the John Templeton Foundation Grant 62840. L.D.G. and D.K. are supported in part by the Simons Investigator Grant No. 144924, and D.K. is supported in part by the U.S. National Science Foundation (NSF) under Grant No. PHY-1818899.
Appendix A Derivation of the electromagnetic tensor for the rotating MBH
@@
To derive the electromagnetic tensor in the rotating magnetically charged spacetime, we follow the complex transformation algorithm [36, 2], starting from the static spherically symmetric MBH described by the Reissner–Nordström metric.
@@
$A_{\mu}=\left(\frac{Pa\cos\theta}{\rho^{2}},0,0,-\frac{r^{2}+a^{2}}{\rho^{2}}P\cos\theta\right).$ (63)
@@
$F_{\mu\nu}=\begin{bmatrix}0&\frac{2Par\cdot\cos\theta}{\rho^{4}}&\frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\\ -\frac{2Par\cdot\cos\theta}{\rho^{4}}&0&0&\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}\\ -\frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0&0&\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}\\ 0&-\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}&-\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\end{bmatrix}$. (64)
As in Section II, we have defined the electromagnetic potential on the northern-hemisphere patch, while the southern patch is related by a gauge transformation.
Appendix B Equations of Motion of Charged Particles Around a Rotating MBH